OpenStudy (asapbleh):

Let r= be an arbitrary vector. In each part, describe the set of all points (x,y) in 2-space that satisfy the stated condition.

OpenStudy (asapbleh):

\[\left| \left| r \right| \right|\le1\]

OpenStudy (dumbcow):

isn't that simply the unit circle

OpenStudy (asapbleh):

How would you do this problem? im stuck

OpenStudy (xapproachesinfinity):

those are the points equidistant from a fixed origin

OpenStudy (asapbleh):

are they outside the radius or inside?

OpenStudy (dumbcow):

inside

OpenStudy (xapproachesinfinity):

yes inside

OpenStudy (asapbleh):

oh because its less than? oh right. duh

OpenStudy (asapbleh):

im having dumb shocks atm

OpenStudy (xapproachesinfinity):

less than or equal! if it is just than that would be another story i think

OpenStudy (asapbleh):

does this apply in 3-D aswell?

OpenStudy (xapproachesinfinity):

well if the origin is fixed you would get a sphere! but if the origin is chosen at random you have something rather different

OpenStudy (xapproachesinfinity):

i was trying to picture this but sucks not good enough at \(R^3\)

OpenStudy (xapproachesinfinity):

@SithsAndGiggles

OpenStudy (asapbleh):

ahh

OpenStudy (xapproachesinfinity):

perhaps siths will give us a nice interpretation here^_^

OpenStudy (asapbleh):

alright, ima do my longass calc hw and look up from time to time at my comp screen

OpenStudy (xapproachesinfinity):

this is calc?

OpenStudy (asapbleh):

isnt it not? im taking calculus 3 rn

OpenStudy (asapbleh):

is*

OpenStudy (xapproachesinfinity):

yes should be^_^

OpenStudy (anonymous):

\[\|r\|=\sqrt{x^2+y^2}\] If \(r=1\) (at most), then you have the circle with radius 1 centered at the origin. For any value of \(r\) less than 1 (but obviously greater than 0), the vector will occupy all points *inside* the circle. So you have all the points on the boundary of the circle as well as the points surrounded by the circle. This region is called a disk (centered at (0,0) with radius 1).

OpenStudy (asapbleh):

so this would apply in 3d realm except it would just be occupying inside, outside, and on the sphere?

OpenStudy (xapproachesinfinity):

let me ask you this question! does the center have to be at (0,0) since it said arbitrary vector?

OpenStudy (asapbleh):

logically, arbitrary vector can be anywhere, but i think they meant to start at the origin becuase it involves 1. iono whenever i see 1 after taking trig, i think of the unit circle

OpenStudy (asapbleh):

i typed out the question exactly how its worded in the textbook.

OpenStudy (xapproachesinfinity):

well we could have a circle with radius at any given center we would still get the circle as long as the center is a fixed point?

OpenStudy (xapproachesinfinity):

radius 1*

OpenStudy (xapproachesinfinity):